# Simultaneous Music Separation and Generation Using Multi-Track Latent Diffusion Models

### v2

@@ I Introduction
Audio source separation refers to the process of isolating individual sound elements from a mixture of sounds. This method is critical in numerous areas, particularly in music production. Recent advances with deep learning models have significantly improved separation quality, utilizing two different approached: discriminative [1, 2, 3, 4, 5, 6, 7] and generative [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Discriminative models focus on directly mapping the input mixture to its separated sources, whereas generative models aim to learn the distribution of individual sources and their combination into a mixture. Diffusion models [18] have lately become a leading generative method in audio source separation tasks [19, 20, 21, 22, 23, 24].
On the other hand, music generation in the domain of raw audio has seen significant advancements with the development of deep learning techniques [25, 26, 27, 28, 29, 30]. These systems create audio content either unconditionally or conditioned on various modalities, often focusing on generating music based on text descriptions of musical genres, mood, and other attributes. Having demonstrated their ability to learn complex data distributions, such as raw audio, diffusion models have had a profound impact on music generation in the audio domain [31, 32, 33, 34]. An alternative paradigm for music generation, which involves using existing musical tracks or melodic hints to generate the remaining music in response to the musical context, has been explored in music-to-music and musical arrangement generation models [35, 36, 37, 38, 39].
@@ II-B Separation and Generation Trade-off
We utilize this paradigm to balance the trade-off between source separation and generation. After simultaneously training the unconditional and conditional models, we vary the guidance scale weight $w$ during inference. By adjusting $w$ between 0 and 1, we effectively switch between source separation and total music generation modes of our model.
@@ II-C Music Separation
To achieve successful music source separation, we introduced a condition over the generation process using the latent representation of the mixture $c=x_{\textit{mix\_latent}}$, resulting in $p_{\theta}(z_{0}|x_{\textit{mix\_latent}})$. The mixture latent $x_{\textit{mix\_latent}}$ is obtained similarly to $x_{\text{latent}}$ through STFT, Mel operations, and a VAE encoder, as shown in Fig. 1. To ensure strong adherence to the conditioning and prevent the model from ”forgetting” the mixture, we implemented a different conditioning method than those used in Audio/MusicLDM models. Instead of using time embedding and attention layers, we add the mixture conditioning directly to every layer of U-Net model. This is accomplished by processing the latent representation of mixture with down-sampling using average pooling, matching the corresponding sizes of the U-Net layers. Additionally, the channel numbers are also matched by repeating the re-sampled mixture as needed. During inference, we apply a CFG weight $w\geq 1$ to ensure strong adherence to the conditioning, enabling effective source separation.
@@ III-A Dataset
III-B Training and Evaluation
The hyperparameter setting for our model largerly follows MusicLDM [31], with extension of the LDM model by modifying the U-Net architecture to handle a 3D latent space with an additional track numbers dimension $S=4$. We used a VAE compression ratio $r=4$, resulting in a latent space shape of $S\times C\times\frac{T}{r}\times\frac{F}{r}=4\times 8\times 256\times 16$, representing stems, channels, time, and frequency, respectively. We used a U-Net channel numbers of 128, 256, 384, and 640 for the encoder and the reversed for the decoder block. Unlike the original Music/AudioLDM, we opted for a simple attention layer instead of a spatial transformer for the U-Net’s attention mechanism, as the latter, designed for 2D image-like data, was incompatible with our model’s 3D latent space. We used pretrained components of MusicLDM—VAE and the HiFi-GAN vocoder—from publicly available checkpoints. We trained our model with unconditional dropout rate of $0.1$, Adam optimizer and learning rate of $3\times 10^{-5}$ for 100 epochs. The total number of steps for DDPM was set at $N=1000$ during training. We used Denoising Diffusion Implicit Models (DDIM) [50] sampler with $N=200$ steps during inference.
For the evaluation of the separation task, we used the Mel mean square distance and a version of the Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) [51] metric applied to flattened Mel-spectrograms. In our scenario, metrics that compare waveforms directly are not applicable because our model uses HiFi-GAN, which obscures direct waveform comparison due to phase reconstruction differences. For generative tasks, we used the Frechet Audio Distance (FAD) [52] metric, a widely recognized benchmark in music evaluation. Following MSDM, in arrangement generation, we used the FAD calculation protocol from [38], where generated tracks are mixed with originals of given tracks and the FAD score is calculated for the resultant mixtures.
TABLE II: Separation performance with SI-SDR (on flattened Mel-spectrogram) and MSE (on Mel-spectrogram) for all tracks (B: Bass, D: Drums, G: Guitar, P: Piano). Our model is compared with our baseline, MSDM.
Si-SDR (on Mel) $\uparrow$ MSE (on Mel)$\downarrow$
Model B D G P B D G P
VAE 35.39 34.25 30.05 32.75 0.28 0.16 0.89 0.54
MSDM [24] 16.75 18.03 16.33 16.37 6.50 3.20 12.16 9.11
MSG-LD (w=1) 19.77 22.23 18.14 18.77 1.86 1.33 5.79 3.07
MSG-LD (w=2) 20.52 22.29 19.02 19.28 1.52 1.01 4.35 2.50
@@ IV-A Source Separation
For the source separation task, we experimented with a CFG weight $w\geq 1.0$ during inference in our model. For baseline comparison, we used the publicly available pretrained model from MSDM and performed separation using the Dirac algorithm, applying the best-performing hyperparameters as reported in their paper. The resulting audio was resampled, and Mel-spectrograms were extracted for direct comparison with our metrics. As reported in Table II, our model significantly outperforms MSDM in the separation task. Although we acknowledge that setting $w>1$ may push the generated samples out of distribution, we empirically observed slightly improved performance with $w=2.0$, which we report in the table. The VAE metrics in the table represent the audio processed through the VAE and compared with the original tracks. Since all generated audio passes through the VAE, these metrics serve as a quality benchmark for our model.

### v3

@@ I Introduction
Audio source separation refers to the process of isolating individual sound elements from a mixture of sounds. This method is critical in numerous areas, particularly in music production. Recent advances with deep learning models have significantly improved separation quality, utilizing two different approached: discriminative [1, 2, 3, 4, 5, 6, 7] and generative [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Discriminative models focus on directly mapping the input mixture to its separated sources, whereas generative models aim to learn the distribution of individual sources and their combination into a mixture. Diffusion models [18] have lately become a leading generative method in audio source separation tasks [19, 20, 21, 22, 23].
On the other hand, music generation in the domain of raw audio has seen significant advancements with the development of deep learning techniques [24, 25, 26, 27, 28, 29]. These systems create audio content either unconditionally or conditioned on various modalities, often focusing on generating music based on text descriptions of musical genres, mood, and other attributes. Having demonstrated their ability to learn complex data distributions, such as raw audio, diffusion models have had a profound impact on music generation in the audio domain [30, 31, 32, 33]. An alternative paradigm for music generation, which involves using existing musical tracks or melodic hints to generate the remaining music in response to the musical context, has been explored in music-to-music and musical arrangement generation models [34, 35, 36, 37, 38, 39].
@@ II-B Separation and Generation Trade-off
We utilize CFG paradigm to balance the trade-off between source separation and generation. After simultaneously training the unconditional and conditional models, we vary the guidance scale weight $w$ during inference. By adjusting $w$ between 0 and 1, we effectively switch between source separation and total music generation modes of our model. Notably, our approach differs from posterior approximation methods like Dirac and Gaussian used in MSDM, which we found incompatible with LDM due to the non-linear relationship between mixtures and sources in the latent space.
@@ II-C Music Separation
To achieve successful music source separation, we introduced a condition over the generation process using the latent representation of the mixture $c=x_{\textit{mix\_latent}}$, resulting in $p_{\theta}(z_{0}|x_{\textit{mix\_latent}})$. The mixture latent $x_{\textit{mix\_latent}}$ is obtained similarly to $x_{\text{latent}}$ through STFT, Mel operations, and a VAE encoder, as shown in Fig. 1. To ensure strong adherence to the conditioning and avoid ”forgetting” the mixture, we replaced FiLM [49] used in MusicLMD with direct mixture conditioning at every layer of the U-Net. This is accomplished by processing the latent representation of mixture with down-sampling using average pooling, matching the corresponding sizes of the U-Net layers. Additionally, the channel numbers are also matched by repeating the re-sampled mixture as needed. During inference, we apply a CFG weight $w\geq 1$ to ensure strong adherence to the conditioning, enabling effective source separation.
@@ III-A Dataset
III-B Model, Training, and Evaluation Setup
The hyperparameter setting for our model largerly follows MusicLDM [30], with extension of the LDM model by modifying the U-Net architecture to handle a 3D latent space with an additional track numbers dimension $S=4$. We used a VAE compression ratio $r=4$, resulting in a latent space shape of $S\times C\times\frac{T}{r}\times\frac{F}{r}=4\times 8\times 256\times 16$, representing stems, channels, time, and frequency, respectively. We used a U-Net channel numbers of 128, 256, 384, and 640 for the encoder and the reversed for the decoder block. Unlike the original Music/AudioLDM, we opted for a simple attention layer instead of a spatial transformer for the U-Net’s attention mechanism, as the latter, designed for 2D image-like data, was incompatible with our model’s 3D latent space. We used pretrained components of MusicLDM—VAE and the HiFi-GAN vocoder—from publicly available checkpoints. We trained our model with unconditional dropout rate of $0.1$, Adam optimizer and learning rate of $3\times 10^{-5}$ for 100 epochs. The total number of steps for DDPM was set at $N=1000$ during training. We used Denoising Diffusion Implicit Models (DDIM) [51] sampler with $N=200$ steps during inference. Our model has  305M parameters in the LDM and  128M in the VAE and vocoder, comparable to baseline models.
We evaluated the separation task using the Mel mean square distance, as SI-SDR [52] is unsuitable for our model. The use of HiFi-GAN introduces phase reconstruction differences, preventing direct waveform comparison. For generative tasks, we used the Frechet Audio Distance (FAD) [53] metric, a widely recognized benchmark in music evaluation. Following MSDM, in arrangement generation, we used the FAD calculation protocol from [37], where generated tracks are mixed with originals of given tracks and the FAD score is calculated for the resultant mixtures.
TABLE II: Separation performance with MSE (on Mel-spectrogram) for all tracks (B: Bass, D: Drums, G: Guitar, P: Piano). Our model is compared with our baseline, MSDM.
MSE (on Mel)$\downarrow$
Model B D G P
VAE 0.28 0.16 0.89 0.54
MSDM [23] 6.50 3.20 12.16 9.11
MSG-LD (w=1) 1.86 1.33 5.79 3.07
MSG-LD (w=2) 1.52 1.01 4.35 2.50
@@ IV-A Source Separation
For the source separation task, we experimented with a CFG weight $w\geq 1.0$ during inference of our model. For baseline comparison, we used the publicly available pretrained model from MSDM and performed separation using the Dirac algorithm, applying the best-performing hyperparameters as reported in their paper. The resulting audio was resampled, and Mel-spectrograms were extracted for direct comparison with our metrics. As reported in Table II, our model significantly outperforms MSDM in the separation task. Although we acknowledge that setting $w>1$ may push the generated samples out of distribution, we empirically observed slightly improved performance with $w=2.0$, which we report in the table. We also tested $w>2.0$ and found that it decreased the quality of the generated samples. The VAE results in the table’s top row benchmark the quality of audio for our model, showing sources processed through the VAE against the original tracks.